# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 3.3. $d$-semistable and generic open Kulikov surfaces
Recall from [72, Definition 2.5] that a stable map $f:C\rightarrow\mathcal{X}_{0}$ is predeformable if $f(C)$ does not contain any stratum of $\mathcal{X}_{0}$, and, for every node $p=C_{i}\cap C_{j}\in C$ with $f(p)$ contained in a component $D_{ij}=V_{i}\cap V_{j}$ of the double locus of $\mathcal{X}_{0}$, we have $f(C_{i})\subset V_{i}$, $f(C_{j})\subset V_{j}$, and the contact order of $f(C_{i})$ along $D_{ij}$ at $f(p)$ is equal to the contact order of $f(C_{j})$ along $D_{ij}$ at $f(p)$.
@@
Let $\mathcal{X}_{0}$ be an open Kulikov surface. A degenerate $(-2)$-curve in $\mathcal{X}_{0}$ is the image $C$ of a predeformable genus zero stable map to $\mathcal{X}_{0}$, such that, denoting by $C_{i}$ the irreducible components of $C$, we have
$\sum_{i}C_{i}^{2}=-2\,,$
where $C_{i}^{2}$ is the self-intersection of $C_{i}$ in the unique irreducible component of $\mathcal{X}_{0}$ containing $C_{i}$.
@@
In Definition 3.19, we consider the cases $\widetilde{\mathcal{X}}_{0}\neq\mathcal{X}_{0}$ for the following reason: if $\mathcal{X}_{0}$ admits a smoothing $\mathcal{X}$, then one wants “generic” to be preserved under taking the central fiber $\widetilde{\mathcal{X}}_{0}$ of a log resolution of a base change of $\mathcal{X}$.
@@
First note that the number of conditions to check in Definition 3.19 is countable and Zariski open in $\mathrm{Hom}(\mathbb{L},\mathbb{C}^{\star})$. Hence, it is enough to check that the locus in $\mathrm{Hom}(\mathbb{L},\mathbb{C}^{\star})$ where each of those conditions is satisfied is non-empty. By [41, Proposition 4.1], a log Calabi–Yau surface $(V_{i},D_{i})$ corresponding to a general point of $\mathrm{Hom}(\mathbb{L}_{i},\mathbb{C}^{\star})$ does not contain any internal $(-2)$-curve.
Now, consider a general degenerate $(-2)$-curve $C=C_{i_{0}}\cup\cdots\cup C_{i_{k}}$ in $\widetilde{\mathcal{X}}_{0}$, as in Definition 3.17. If $C$ is the pre-image of a degenerate $(-2)$-curve in $\mathcal{X}_{0}$, then one can assume $\widetilde{\mathcal{X}}_{0}=\mathcal{X}_{0}$, and there exists an open Kulikov surface $\mathcal{X}_{0}^{\prime}$, obtained from $\mathcal{X}_{0}$ by a sequence of M1 modifications, and containing an internal $(-2)$-curve. Moreover, it follows from [7, §7C] that $\mathrm{Hom}(\mathbb{L},\mathbb{C}^{\star})$ for $\mathcal{X}_{0}$ and $\mathcal{X}_{0}^{\prime}$ are naturally identified. Hence, the result follows in this case from the previous argument for internal $(-2)$-curves. If $C$ is not the pre-image of a degenerate $(-2)$-curve in $\mathcal{X}_{0}$, then there is at least one component $D_{ij}$ of the double locus of $\widetilde{\mathcal{X}}_{0}$ which intersects $C$ and is contracted in $\mathcal{X}_{0}$. On the other hand, the existence of $C$ imposes a non-trivial condition on the positions of the internal exceptional curves corresponding to the univalent vertices of the dual complex of $C$, and so defines a closed proper subset of $\prod_{k}\mathrm{Hom}(\mathbb{L}_{k},\mathbb{C}^{\star})$. If we generically change those positions, so that $C$ disappears, then it follows from the proof of Theorem 3.16 that one can make $\mathcal{X}_{0}$ $d$-semistable by changing the gluing along the double locus of $\mathcal{X}_{0}$. As $D_{ij}$ is contracted in $\mathcal{X}_{0}$, the gluing along $D_{ij}$ in $\widetilde{\mathcal{X}}_{0}$ is not affected, and so the class of $C$ remains non-effective. ∎
@@ 3.4.1. Log smooth deformation theory
Recall that on a scheme $X$, a log structure is given by a sheaf of commutative monoids $\mathcal{M}_{X}$, together with a homomorphism of sheaves $\alpha_{X}:\mathcal{M}_{X}\to\mathcal{O}_{X}$ between multiplicative monoids, inducing an isomorphism $\alpha_{X}^{-1}(\mathcal{O}_{X}^{\times})\rightarrow\mathcal{O}_{X}^{\times}$. This isomorphism allows us to identify $\mathcal{O}_{X}^{\times}$ as a subsheaf of $\mathcal{M}_{X}$. We call a scheme $X$ endowed with a log structure a log scheme, and denote it by $X^{\dagger}:=(X,\mathcal{M}_{X})$ – we suppress the $\alpha_{X}$ from the notation, at situations it is clear what is is. Throughout this paper we assume that all log schemes are fine and saturated [78, I,§1.3]. The sheaf ¯$\overline{\mathcal{M}}_{X}:=\mathcal{M}_{X}/\mathcal{O}_{X}^{\times}$ known as the ghost sheaf or characteristic sheaf, captures the key combinatorial information about the log structure and in particular, the tropicalization $\Sigma(X)$ of $X$, which is an abstract polyhedral cone complex – see [1], §2.1 for details. In brief, the tropicalization $\Sigma(X)$ is given by a collection of cones along with face maps between them. For every geometric point $\bar{x}\rightarrow X$, there is one cone $\sigma_{\bar{x}}:=\mathrm{Hom}(\overline{\mathcal{M}}_{X,\bar{x}},\mathbb{R}_{\geq 0})$, and if $\bar{x}$ specializes to $\bar{y}$, there is a generization map ¯$\overline{\mathcal{M}}_{X,\bar{y}}\rightarrow\overline{\mathcal{M}}_{X,\bar{x}}$. This leads dually to a map $\sigma_{\bar{x}}\rightarrow\sigma_{\bar{y}}$, and the assumption that our log structures implies this is an inclusion of faces.
@@
Given a log scheme $Y^{\dagger}=(Y,\mathcal{M}_{Y})$ and a morphism of schemes $f:X\to Y$, one can define an induced log structure, also known as pull-back log structure on $X$, which we denote by $f^{*}\mathcal{M}_{Y}$ – see [13, Defn.2.16]. Pulling-back the divisorial log structure on a toric variety $X=\mathrm{Spec}\,\mathbb{C}[Q]$ to the to a torus fixed point $\mathrm{Spec}\mathbb{C}$ (assuming $Q^{\times}=\{0\}$, so that such a point exists), we obtain the standard log structure on $\mathrm{Spec}\,\mathbb{C}$ given by
@@
Throughout this paper, unless stated otherwise, we assume $Q:=\mathbb{N}$ and denote the standard log point by
${\mathrm{Spec}\,\mathbb{C}}^{\dagger}:=(\mathrm{Spec}\,\mathbb{C},\mathbb{N}\oplus\mathbb{C}^{\times})\,.$
@@
Similarly, as for scheme theoretic morphisms we have the notion of smoothness which can be defined using an infinitesimal lifting criterion, for log morphisms, we have the analog of this notion, given by log smoothness. We call a morphism log smooth, if and only if it is locally toroidal, in the sense of [56, (3.5)]. In the following sections, we study the deformation theory of the normal crossing surface $\mathcal{X}_{0}$, which satisfies “log smoothness”, since by Theorem 3.21 it is $d$-semistable, and as shown in [61] and in [56, Theorem 11.2], this implies log smoothness. While log smooth deformation theory is developed in a general set-up by Fumiharu Kato [56], it suffices for our purposes to focus into the log deformations in the normal crossing crossing situation, following [61]. We do so in the following section.
@@ 3.4.2. Smoothing of $\mathcal{X}_{0}$
Throughout this section, for every log morphism $f^{\dagger}:X^{\dagger}\rightarrow Y^{\dagger}$, between two log schemes $X$ and $Y$ we denote by $T_{X^{\dagger}/Y^{\dagger}}$ (resp. $\Omega_{X^{\dagger}/Y^{\dagger}}$) the associated relative log tangent (resp. cotangent) sheaf, which reduces to a vector bundle in the situation when the morphism $f^{\dagger}$ is log smooth [57, Proposition 3.10].
@@
to the standard log point, whose tropicalization is the height function
$h\colon(C(P),C(\mathscr{P}))\to(\mathbb{R}_{\geq 0},\Sigma_{\mathbb{A}^{1}})\,,$
given by projection onto the second factor.
@@
Since $\mathcal{X}_{0}$ is $d$-semistable, by [61, Proposition 1.1] it admits a log structure which is log smooth over the standard log point. Since by construction $(P,\mathscr{P})$ is the dual intersection complex of $\mathcal{X}_{0}$, the statement about the tropicalization is immediate by the description of the tropicalization of a map, as explained in §3.4.1. ∎
The key ingredient to show the existence of the smoothing of $\mathcal{X}_{0}$, is the “vanishing result” shown in Lemma 3.25. Before that, we first need the following observation for smooth log Calabi–Yau surfaces.
@@
then $a_{i}=0$ for all $3\leq i\leq n$. First, taking the intersection with $D_{2}$ of both sides of Equation (3.4), we get $a_{3}(D_{2}\cdot D_{3})+\ldots a_{n}(D_{2}\cdot D_{n})=0$. Since $D_{2}\cdot D_{3}=1$ and $D_{2}\cdot D_{i}=0$ for all $3\leq i\leq n$, as $D_{1}+\ldots+D_{n}$ is cyclically ordered, this gives us $a_{3}=0$. Thus, Equation 3.4 reduces to $a_{4}D_{4}+\ldots a_{n}D_{n}=0$. Now, taking the intersection with $D_{3}$ of both sides, analogously as before, we get $a_{4}=0$. Proceeding similarly, eventually we get $a_{i}=0$ for all $3\leq i\leq n$. ∎
@@
We adapt to our situation the argument given in the proof of [33, Lemma 5.9] (where the log cotangent bundle is denoted by $\Lambda^{1}_{X}$) in the context of degenerations of K3 surfaces – see also [83, Theorem 5.2] for a similar result in the context of degenerations of Enriques surfaces.
As in [33, Corollary 3.6] or [61, Proof of Theorem 4.2], it follows from the definition of the log cotangent bundle that we have a short exact sequence
@@
To prove Lemma 3.24, we have to show that
$H^{0}(\mathcal{Z},(\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}})(-\partial\mathcal{Z}))=0\,,$
that is, by (3.8), that the map $\delta$ is injective. Explicitly, an element of $H$ is a sequence $(a_{e})_{e\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial Z}}$ of numbers $a_{e}\in\mathbb{C}$, satisfying the balancing condition $a_{e_{1}}+a_{e_{2}}+a_{e_{3}}=0$ each time we have three edges $e_{1},e_{2},e_{3}\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial Z}$ bounding a triangle $f\in\mathscr{P}^{[2]}_{\mathcal{Z}}$. Let $(a_{e})_{e\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial Z}}$ be in the kernel of $\delta$, that is, that the image of $\sum_{e\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}}a_{e}[X^{e}]$ in $H^{2}(X^{v},\mathbb{C})$ is zero for every $v\in\mathscr{P}^{[1]}_{\mathcal{Z}}$. For every vertex $v\in\mathscr{P}^{[0]}$, define the distance to the boundary $d(v)$ as $d(v)=0$ if $v\in\partial P$, and else as the smallest integer such that there exists a sequence $(e_{1},\dots,e_{d(v)})$ of edges $e_{i}\in\mathscr{P}^{[1]}$ such that $e_{1}$ is adjacent to $v$, $e_{i+1}$ is adjacent to $e_{i}$, and $e_{d(v)}\in\partial P$. We will prove that $a_{e}=0$ for all $e\in\mathscr{P}^{[1]}_{\mathcal{Z}}$, by showing by induction on $d(v)\geq 0$ that, for every vertex $v\in\mathscr{P}^{[0]}$, we have $a_{e}=0$ for all edges $e\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}$ adjacent to $v$.
@@
Recall that the affinization map $f_{0}:\mathcal{X}_{0}\rightarrow\overline{\mathcal{X}}_{0}$ is a projective surjective birational morphism. Since $\overline{\mathcal{X}}_{0}$ is affine, then, by [50, Proposition III.8.5] (a degenerate case of the Leray spectral sequence), we have
@@
Fibers of $f_{0}$ over points in $\overline{\mathcal{X}}_{0}$ distinct from the central point $x_{0}$ are of dimension at most one, and so, by [50, Corollary III.11.2], the sheaf $R^{2}(f_{0})_{\star}\,T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}$ is supported on the central point $x_{0}$. Hence, it is enough to show that its stalk at $x_{0}$ is 00:
@@
As the sheaf $R^{2}(f_{0})_{\star}\,T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}$ is supported $x_{0}$, the formal completion of the stalk at $x_{0}$ is equal to the stalk at $x_{0}$ itself, and so, by the theorem on formal functions [50, Theorem 11.1], we have
@@
The irreducible components of $\mathcal{X}_{0}^{(0)}$ are the proper surfaces $X^{v}_{\mathscr{P}}$, indexed by the interior vertices $v\in\mathscr{P}^{[0]}$, and the proper curves $X^{e}_{\mathscr{P}}$ indexed by the set $\mathscr{P}^{[1]}_{\mathcal{E}}$ of interior edges $e\in\mathscr{P}^{[1]}$ connecting pairs of boundary vertices. We denote by
$\mathcal{X}^{(0)}_{0}=\mathcal{Z}\cup\mathcal{E}$
the corresponding decomposition, where
@@
By Lemma 3.8, the dualizing sheaf $\omega_{\mathscr{X}_{0}}$ is trivial, and so, by adjunction, the dualizing sheaf $\omega_{\mathcal{Z}}$ of $\mathcal{Z}$ is given by the line bundle
$\omega_{\mathcal{Z}}=\mathcal{O}_{\mathcal{Z}}(-\partial\mathcal{Z})\,.$
Hence, by Serre duality on $\mathcal{Z}$, and using that $T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}}$ is locally free on $\mathcal{Z}$, we obtain
@@
Finally, we are ready to show the existence of a smoothing of $\mathcal{X}_{0}$.
@@
Since the obstruction group $H^{2}(\mathcal{X}_{0},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\mathbb{C}^{\dagger}})$ vanishes by Lemma 3.25, the result follows by log smooth deformation theory as in [61, Corollary 2.4]. ∎

### 30 Mar 2024 v2

@@ 3.3. $d$-semistable and generic open Kulikov surfaces
Recall from [73, Definition 2.5] that a stable map $f:C\rightarrow\mathcal{X}_{0}$ is predeformable if $f(C)$ does not contain any stratum of $\mathcal{X}_{0}$, and, for every node $p=C_{i}\cap C_{j}\in C$ with $f(p)$ contained in a component $D_{ij}=V_{i}\cap V_{j}$ of the double locus of $\mathcal{X}_{0}$, we have $f(C_{i})\subset V_{i}$, $f(C_{j})\subset V_{j}$, and the contact order of $f(C_{i})$ along $D_{ij}$ at $f(p)$ is equal to the contact order of $f(C_{j})$ along $D_{ij}$ at $f(p)$.
@@
Let $\mathcal{X}_{0}$ be an open Kulikov surface. A degenerate $(-2)$-curve $C$ in $\mathcal{X}_{0}$ is the image of a predeformable genus zero stable map to $\mathcal{X}_{0}$, such that $\sum_{i}C_{i}^{2}=-2$, where $C_{i}$ are the irreducible components of $C$, and $C_{i}^{2}$ is the self-intersection of $C_{i}$ in the unique irreducible component of $\mathcal{X}_{0}$ containing $C_{i}$.
@@
If $\mathcal{X}_{0}$ admits a smoothing $\mathcal{X}$, then the “generic” condition is preserved under replacing $\mathcal{X}_{0}$ by the central fiber $\widetilde{\mathcal{X}}_{0}$ of a log resolution of a base change of $\mathcal{X}$.
@@
First note that the number of conditions to check in Definition 3.19 is countable and that each of them is satisfied on a Zariski open subset in $\mathrm{Hom}(\mathbb{L},\mathbb{C}^{\star})$. Hence, it is enough to check that the locus in $\mathrm{Hom}(\mathbb{L},\mathbb{C}^{\star})$ where each of those conditions is satisfied is non-empty. By [42, Proposition 4.1], a log Calabi–Yau surface $(V_{i},D_{i})$ corresponding to a general point of $\mathrm{Hom}(\mathbb{L}_{i},\mathbb{C}^{\star})$ does not contain any internal $(-2)$-curve.
Now, consider a general degenerate $(-2)$-curve $C=C_{i_{0}}\cup\cdots\cup C_{i_{k}}$ in $\widetilde{\mathcal{X}}_{0}$, as in Definition 3.17. If $C$ is the pre-image of a degenerate $(-2)$-curve in $\mathcal{X}_{0}$, then we can assume that $\widetilde{\mathcal{X}}_{0}=\mathcal{X}_{0}$, and there exists an open Kulikov surface $\mathcal{X}_{0}^{\prime}$, obtained from $\mathcal{X}_{0}$ by a sequence of M1 modifications, and containing an internal $(-2)$-curve. Moreover, it follows from [7, §7C] that $\mathrm{Hom}(\mathbb{L},\mathbb{C}^{\star})$ for $\mathcal{X}_{0}$ and $\mathcal{X}_{0}^{\prime}$ are naturally identified. Hence, the result follows in this case from the previous argument for internal $(-2)$-curves. If $C$ is not the pre-image of a degenerate $(-2)$-curve in $\mathcal{X}_{0}$, then there is at least one component $D_{ij}$ of the double locus of $\widetilde{\mathcal{X}}_{0}$ which intersects $C$ and is contracted in $\mathcal{X}_{0}$. On the other hand, the existence of $C$ imposes a non-trivial condition on the positions of the internal exceptional curves corresponding to the univalent vertices of the dual complex of $C$, and so defines a closed proper subset of $\prod_{k}\mathrm{Hom}(\mathbb{L}_{k},\mathbb{C}^{\star})$. If we generically change those positions, so that $C$ disappears, then it follows from the proof of Theorem 3.16 that one can make $\mathcal{X}_{0}$ $d$-semistable by changing the gluing along the double locus of $\mathcal{X}_{0}$. As $D_{ij}$ is contracted in $\mathcal{X}_{0}$, the gluing along $D_{ij}$ in $\widetilde{\mathcal{X}}_{0}$ is not affected, and so the class of $C$ remains non-effective. ∎
@@ 3.4.1. Log smooth deformation theory
Recall that on a scheme $X$, a log structure is given by a sheaf of commutative monoids $\mathcal{M}_{X}$, together with a homomorphism of sheaves $\alpha_{X}:\mathcal{M}_{X}\to\mathcal{O}_{X}$ between multiplicative monoids, inducing an isomorphism $\alpha_{X}^{-1}(\mathcal{O}_{X}^{\times})\rightarrow\mathcal{O}_{X}^{\times}$. This isomorphism allows us to identify $\mathcal{O}_{X}^{\times}$ as a subsheaf of $\mathcal{M}_{X}$. We call a scheme $X$ endowed with a log structure a log scheme, and denote it by $X^{\dagger}:=(X,\mathcal{M}_{X})$ – we suppress the $\alpha_{X}$ from the notation, at situations it is clear what is is. Throughout this paper we assume that all log schemes are fine and saturated [79, I,§1.3]. The sheaf ¯$\overline{\mathcal{M}}_{X}:=\mathcal{M}_{X}/\mathcal{O}_{X}^{\times}$ known as the ghost sheaf or characteristic sheaf, captures the key combinatorial information about the log structure and in particular, the tropicalization $\Sigma(X)$ of $X$, which is an abstract polyhedral cone complex – see [1], §2.1 for details. In brief, the tropicalization $\Sigma(X)$ is given by a collection of cones along with face maps between them. For every geometric point $\bar{x}\rightarrow X$, there is one cone $\sigma_{\bar{x}}:=\mathrm{Hom}(\overline{\mathcal{M}}_{X,\bar{x}},\mathbb{R}_{\geq 0})$, and if $\bar{x}$ specializes to $\bar{y}$, there is a generization map ¯$\overline{\mathcal{M}}_{X,\bar{y}}\rightarrow\overline{\mathcal{M}}_{X,\bar{x}}$. This leads dually to a map $\sigma_{\bar{x}}\rightarrow\sigma_{\bar{y}}$, and the assumption that our log structures implies this is an inclusion of faces.
@@
Given a log scheme $Y^{\dagger}=(Y,\mathcal{M}_{Y})$ and a morphism of schemes $f:X\to Y$, one can define an induced log structure, also known as pull-back log structure on $X$, which we denote by $f^{*}\mathcal{M}_{Y}$ – see [13, Defn.2.16]. For a monoid $Q$ such that $Q^{\times}=\{0\}$, pulling-back the divisorial log structure on the affine toric variety $X=\mathrm{Spec}\,\mathbb{C}[Q]$ to its torus fixed point $\mathrm{Spec}\,\mathbb{C}$, we obtain the standard log structure on $\mathrm{Spec}\,\mathbb{C}$ given by
@@
Throughout this paper, unless stated otherwise, we assume $Q:=\mathbb{N}$ and denote the standard log point by ${\mathrm{Spec}\,\mathbb{C}}^{\dagger}:=(\mathrm{Spec}\,\mathbb{C},\mathbb{N}\oplus\mathbb{C}^{\times})$.
@@
Similarly, as for scheme theoretic morphisms we have the notion of smoothness which can be defined using an infinitesimal lifting criterion, for log morphisms, we have the analog of this notion, given by log smoothness. We call a morphism log smooth, if and only if it is locally toroidal, in the sense of [57, (3.5)]. In the following sections, we study the deformation theory of the normal crossing surface $\mathcal{X}_{0}$, which satisfies “log smoothness”, since by Theorem 3.21 it is $d$-semistable, and as shown in [62] and in [57, Theorem 11.2], this implies log smoothness. While log smooth deformation theory is developed in a general set-up by Fumiharu Kato [57], it suffices for our purposes to consider the normal crossing situation, following [62]. We do so in the following section.
@@ 3.4.2. Smoothing of $\mathcal{X}_{0}$
Throughout this section, for every log morphism $f^{\dagger}:X^{\dagger}\rightarrow Y^{\dagger}$, between two log schemes $X$ and $Y$ we denote by $T_{X^{\dagger}/Y^{\dagger}}$ (resp. $\Omega_{X^{\dagger}/Y^{\dagger}}$) the associated relative log tangent (resp. cotangent) sheaf, which reduces to a vector bundle in the situation when the morphism $f^{\dagger}$ is log smooth [58, Proposition 3.10].
@@
to the standard log point, whose tropicalization is the height function $h\colon(C(P),C(\mathscr{P}))\to(\mathbb{R}_{\geq 0},\Sigma_{\mathbb{A}^{1}})$, given by projection onto the second factor.
@@
Since $\mathcal{X}_{0}$ is $d$-semistable, by [62, Proposition 1.1] it admits a log structure which is log smooth over the standard log point. Since by construction $(P,\mathscr{P})$ is the dual intersection complex of $\mathcal{X}_{0}$, the statement about the tropicalization is immediate by the description of the tropicalization of a map, as explained in §3.4.1. ∎
The key ingredient to show the existence of the smoothing of $\mathcal{X}_{0}$, is the “vanishing result” shown in Lemma 3.25. Its proof uses the following observation for smooth log Calabi–Yau surfaces.
@@
then $a_{i}=0$ for all $3\leq i\leq n$. First, taking the intersection with $D_{2}$ of both sides of Equation (3.3), we get $a_{3}(D_{2}\cdot D_{3})+\ldots a_{n}(D_{2}\cdot D_{n})=0$. Since $D_{2}\cdot D_{3}=1$ and $D_{2}\cdot D_{i}=0$ for all $3\leq i\leq n$, as $D_{1}+\ldots+D_{n}$ is cyclically ordered, this gives us $a_{3}=0$. Thus, Equation (3.3) reduces to $a_{4}D_{4}+\ldots a_{n}D_{n}=0$. Now, taking the intersection with $D_{3}$ of both sides, analogously as before, we get $a_{4}=0$. Proceeding similarly, eventually we get $a_{i}=0$ for all $3\leq i\leq n$. ∎
@@
We adapt to our situation the argument given in the proof of [33, Lemma 5.9] (where the log cotangent bundle is denoted by $\Lambda^{1}_{X}$) in the context of degenerations of K3 surfaces – see also [84, Theorem 5.2] for a similar result in the context of degenerations of Enriques surfaces.
As in [33, Corollary 3.6] or [62, Proof of Theorem 4.2], it follows from the definition of the log cotangent bundle that we have a short exact sequence
@@
To prove Lemma 3.24, we have to show that $H^{0}(\mathcal{Z},(\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}})(-\partial\mathcal{Z}))=0$, that is, by (3.7), that the map $\delta$ is injective. Explicitly, an element of $H$ is a sequence $(a_{e})_{e\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial Z}}$ of numbers $a_{e}\in\mathbb{C}$, satisfying the balancing condition $a_{e_{1}}+a_{e_{2}}+a_{e_{3}}=0$ each time we have three edges $e_{1},e_{2},e_{3}\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial Z}$ bounding a triangle $f\in\mathscr{P}^{[2]}_{\mathcal{Z}}$. Let $(a_{e})_{e\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial Z}}$ be in the kernel of $\delta$, that is, that the image of $\sum_{e\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}}a_{e}[X^{e}]$ in $H^{2}(X^{v},\mathbb{C})$ is zero for every $v\in\mathscr{P}^{[1]}_{\mathcal{Z}}$. For every vertex $v\in\mathscr{P}^{[0]}$, define the distance to the boundary $d(v)$ as $d(v)=0$ if $v\in\partial P$, and else as the smallest integer such that there exists a sequence $(e_{1},\dots,e_{d(v)})$ of edges $e_{i}\in\mathscr{P}^{[1]}$ such that $e_{1}$ is adjacent to $v$, $e_{i+1}$ is adjacent to $e_{i}$, and $e_{d(v)}\in\partial P$. We will prove that $a_{e}=0$ for all $e\in\mathscr{P}^{[1]}_{\mathcal{Z}}$, by showing by induction on $d(v)\geq 0$ that, for every vertex $v\in\mathscr{P}^{[0]}$, we have $a_{e}=0$ for all edges $e\in\mathscr{P}^{[1]}_{\mathcal{Z}\setminus\partial\mathcal{Z}}$ adjacent to $v$.
@@
Recall that the affinization map $f_{0}:\mathcal{X}_{0}\rightarrow\overline{\mathcal{X}}_{0}$ is a projective surjective birational morphism. Since $\overline{\mathcal{X}}_{0}$ is affine, then, by [51, Proposition III.8.5] (a degenerate case of the Leray spectral sequence), we have
@@
Fibers of $f_{0}$ over points in $\overline{\mathcal{X}}_{0}$ distinct from the central point $x_{0}$ are of dimension at most one, and so, by [51, Corollary III.11.2], the sheaf $R^{2}(f_{0})_{\star}\,T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}$ is supported on the central point $x_{0}$. Hence, it is enough to show that its stalk at $x_{0}$ is 00:
@@
As the sheaf $R^{2}(f_{0})_{\star}\,T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}$ is supported $x_{0}$, the formal completion of the stalk at $x_{0}$ is equal to the stalk at $x_{0}$ itself, and so, by the theorem on formal functions [51, Theorem 11.1], we have
@@
The irreducible components of $\mathcal{X}_{0}^{(0)}$ are the proper surfaces $X^{v}_{\mathscr{P}}$, indexed by the interior vertices $v\in\mathscr{P}^{[0]}$, and the proper curves $X^{e}_{\mathscr{P}}$ indexed by the set $\mathscr{P}^{[1]}_{\mathcal{E}}$ of interior edges $e\in\mathscr{P}^{[1]}$ connecting pairs of boundary vertices. We denote by $\mathcal{X}^{(0)}_{0}=\mathcal{Z}\cup\mathcal{E}$ the corresponding decomposition, where
@@
By Lemma 3.8, the dualizing sheaf $\omega_{\mathscr{X}_{0}}$ is trivial, and so, by adjunction, the dualizing sheaf $\omega_{\mathcal{Z}}$ of $\mathcal{Z}$ is given by the line bundle $\omega_{\mathcal{Z}}=\mathcal{O}_{\mathcal{Z}}(-\partial\mathcal{Z})$. Hence, by Serre duality on $\mathcal{Z}$, and using that $T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}}$ is locally free on $\mathcal{Z}$, we obtain
@@
Finally, we are ready to state the existence of a smoothing of $\mathcal{X}_{0}$.
@@
Since the obstruction group $H^{2}(\mathcal{X}_{0},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\mathbb{C}^{\dagger}})$ vanishes by Lemma 3.25, the result follows by log smooth deformation theory as in [62, Corollary 2.4]. ∎